OpenStudy (anonymous):

Solve e^4x=5.7 for x correct to four decimal places.

OpenStudy (anonymous):

calculator exercise for sure start with \[4x=\ln(5.7)\] then compute \[\ln(5.7)\div 4\]

OpenStudy (anonymous):

\[e^A=B\iff A=\ln(B)\] is what you use

Miracrown (miracrown):

We will have to use logarithms. We want to use logs to take the 4x out of the exponent.

OpenStudy (anonymous):

you can do it here if you like, without using a calculator http://www.wolframalpha.com/input/?i=ln%285.7%29%2F4

Miracrown (miracrown):

Any ideas on how we can use logs to accomplish that? @ejune420

OpenStudy (anonymous):

sorry guys my computer froze up. thanks @satellite73 i got it now!